what is the percentage for 36 of 20%
Answer by checkley75(3666)

IF YOU ARE ASKING WHAT IS 20% OF 36(OF GENERALLY MEANS MULTIPLY)
(20%=20/100=.20) THEN .20*36=7.2
IF YOU ARE ASKING 36 IS 20% OF WHAT NUMBER THEN .20X=36 X=36/.20 X=180
-----------------------------------------------------------------------------
20% OUT OF 36 IS .20*36=7.2
---------------------------------------------------------------------------
THINK OF A SIMILAR PROBLEM:
WHAT IS 50% OF 10. 50%=.50 SO .50*10=5
SO 50% OR 1/2 OR .50 OF A NUMBER IS 1/2*40=20, .50*56=29, 50%*90=45
----------------------------------------------------------------------------
HOPE THIS HELPS EXPLAIN HOW YOU GET PERCENTAGES OF A NUMBER. : )